# OpenStereo: A Comprehensive Benchmark for Stereo Matching and Strong Baseline

### v6

@@ Abstract
Stereo matching aims to estimate the disparity between matching pixels in a stereo image pair, which is important to robotics, autonomous driving, and other computer vision tasks. Despite the development of numerous impressive methods in recent years, determining the most suitable architecture for practical application remains challenging. Addressing this gap, our paper introduces a comprehensive benchmark focusing on practical applicability rather than solely on individual models for optimized performance. Specifically, we develop a flexible and efficient stereo matching codebase, called OpenStereo. OpenStereo includes training and inference codes of more than 10 network models, making it, to our knowledge, the most complete stereo matching toolbox available. Based on OpenStereo, we conducted experiments and have achieved or surpassed the performance metrics reported in the original paper. Additionally, we conduct an exhaustive analysis and deconstruction of recent developments in stereo matching through comprehensive ablative experiments. These investigations inspired the creation of StereoBase, a strong baseline model. Our StereoBase ranks 1st on SceneFlow, KITTI 2015, 2012 (Reflective) among published methods and performs best across all metrics. In addition, StereoBase has strong cross-dataset generalization. Code is available at https://github.com/XiandaGuo/OpenStereo.
@@ 1 Introduction
Traditional stereo-matching algorithms typically rely on matching corresponding image regions between the left and right views based on their similarity measures. Several techniques have been proposed in the literature for stereo matching, including methods based on gray-level information [3, 29, 68], region-based approaches [73, 39], and energy optimization methods [44, 21]. With the support of large synthetic datasets [35, 43, 45, 14, 36, 66], CNNs-based stereo matching methods [24, 6, 70, 18, 61, 33, 10, 52, 60, 53, 8, 30, 32] has achieved impressive results. As shown in Figure 1, based on the network pipeline of stereo matching, CNNs-based stereo matching methods can be roughly grouped into two categories [52], including the encoder-decoder network with 2D convolution (ED-Conv2D) [35, 40, 65, 41, 51, 52, 61, 49, 31, 33, 46, 28, 56, 30, 16] and the cost volume matching with 3D convolution (CVM-Conv3D) [24, 6, 67, 55, 70, 18, 71, 12, 74, 15, 1, 10, 2, 47, 59, 60, 54, 9, 62].
However, we find that different studies often employ varying data augmentation strategies, learning rates, learning rate optimization methods, and backbone architectures. This inconsistency makes it difficult to evaluate and compare various methods’ performance accurately. This inconsistency in experimental setups and methodologies makes it difficult to derive conclusive insights and hampers the objective assessment of advancements in stereo matching. Without a standardized benchmark, researchers struggle to identify the true impact of individual components and innovations. There is a lack of clear conclusions and exploration regarding data augmentation strategies, backbone selection, and cost construction methods in stereo matching.
Moreover, not all datasets are accompanied by official evaluation tools. For example, the DrivingStereo [66] dataset does not provide specific evaluation scripts, making comparative assessments challenging. The SceneFlow [35] dataset, with its finalpass and cleanpass data varieties, complicates fair model comparisons. Generalization experiments for stereo matching algorithms typically train on the SceneFlow dataset and evaluate on KITTI2012 [14], KITTI2015 [36], ETH3D [45], and Middlebury [43]. However, due to the absence of a standard protocol for generalization experiments, different papers may yield inconsistent results for the same method. For instance, discussions on the generalization performance of IGEV [60] across works [60, 53, 17] exemplify this issue.
@@
Hence, there’s a pressing need for a comprehensive benchmark study within the stereo-matching community to enhance practicality and ensure consistent comparisons. To achieve this objective, we introduce a versatile stereo-matching codebase: OpenStereo. To promote scalability and adaptability, OpenStereo offers the following features: (1) Modual design, researchers can define a new model without the need to alter the model code itself by simply modifying a yaml configuration file. (2)Various frameworks, including Concatenation-based [6, 70], Correlation-based [61, 51, 52, 2], Interlaced-based [46], Group-wise-correlation-based [60], Combine-based methods [18], and Difference-based [26]. (3)Various datasets, including SceneFlow [35], KITTI2012 [14], KITTI2015 [36], Middlebury [43], ETH3D [45] and DrivingStereo [66] dataset. (4) State-of-the-art methods, including PSMNet [6], GwcNet [18], AANet [61], FADNet++ [52], CFNet [47], STTR [31], CoEx [2], CascadeStereo [15], MobileStereoNet [46] and IGEV [60].
Leveraging OpenStereo, we rigorously reassess various officially stated conclusions by meticulously re-implementing the ablation studies, including data augmentation, backbone architectures, cost construction, disparity regression, and refinement processes. Based on the insights gleaned from these ablation experiments, we introduce StereoBase, a model that sets a new benchmark, surpassing recently proposed methods in terms of performance. StereoBase is powerful and serves as an empirically robust baseline model for stereo matching, demonstrating exceptional efficacy and resilience across diverse testing scenarios.
Our contributions are summarized as follows:
@@ 2.1 Stereo Matching
Stereo Matching with CVM-Conv3D The CVM-Conv3D methods are proposed to improve the performance of depth estimation [24, 6, 67, 55, 70, 18, 71, 12, 74, 15, 1, 10, 2, 47, 59, 60, 54, 9, 62]. These methods learn disparities from a 4D cost volume, mainly constructed by concatenating left feature maps with their corresponding right counterparts across each disparity level [6]. GCNet [24] firstly introduced a novel approach that combines 3D encoder-decoder architecture with a 2D convolutional network to obtain a dense feature representation, which is used to regularize a 4D concatenation volume. Following GCNet [24], PSMNet [6] proposes an approach for regularizing the concatenation volume by leveraging a stacked hourglass 3D convolutional neural network in tandem with intermediate supervision. To enhance the expressiveness of the cost volume and ultimately improve performance in ambiguous regions, GwcNet [70] proposes the group-wise correlation volume and ACVNet [59] proposes the attention concatenation volume. GANet [70] proposes a semi-global aggregation layer and local guided aggregation layer to capture local and the whole-image cost dependencies respectively. Based on GANet [70], DSMNet [71] constructs a robust “domain normalization” method to overcome the challenges in cross-domain generalization. CoEx [2] proposes a novel approach called Guided Cost volume Excitation (GCE), which leverages image guidance to construct a simple channel excitation of the cost volume. IGEV-Stereo [60] leverages an iterative geometry encoding volume to capture local and non-local geometry information, outperforming existing methods on KITTI benchmarks and achieving cross-dataset generalization and high inference efficiency. Besides, as a potential solution to reduce the need for humans to design a good volumetric method, Neural Architecture Search (NAS) [20, 34, 4] has been employed in stereo matching [10].
@@
Stereo Matching with ED-Conv2D The ED-Conv2D methods [35, 40, 65, 41, 51, 52, 61, 49, 31, 33, 46, 28, 56, 30, 16], which adopt networks with 2D convolutions to predict disparity, has been driven by the need for improved accuracy, computational efficiency, and real-time performance. One of the early deep learning-based stereo matching methods, MC-CNN (Matching Cost CNN) [69], was proposed to learn a matching cost function for improving performance in the cost aggregation and optimization stages. Subsequently, Mayer et al [35] present end-to-end networks for the estimation of disparity, called DispNet, which is pure 2D CNN architectures. However, the model still faces challenges in capturing the matching features, resulting in poor estimation results. To overcome this challenge, the correlation layer is introduced in the end-to-end architecture [35, 11, 22, 23] to better capture the relationship between the left and right images. By incorporating this layer, the accuracy of the model is significantly improved. Furthermore, FADNet++ [52] proposes an innovative approach to efficient disparity refinement using residual learning [19] in a coarse-to-fine manner. AutoDispNet [41] applied neural architecture search to automatically design stereo matching network structures. More recently, Croco-Stereo [56] shows that large-scale pre-training can be successful for stereo matching through well-adapted pretext tasks. This method can achieve state-of-the-art performance without using task-specific designs, like correlation volume or iterative estimation.
@@ 3 OpenStereo
In recent years, there has been a proliferation of new frameworks and evaluation datasets for stereo matching. However, the lack of a unified and fair evaluation platform in this field is a significant issue that cannot be ignored. To address this challenge and promote academic research and practical application we have developed OpenStereo, a pyTorch-based [38] toolbox that provides a reliable and standardized evaluation framework for stereo matching.
Figure 2: The design principles of proposed codebase OpenStereo.
@@ 3.1 Design Principles of OpenStereo
Compatibility with various frameworks. Currently, more and more stereo matching methods have emerged, such as Concatenation-based [24, 6, 70, 10], Correlation-based [61, 51, 52, 2], Group-wise-correlation-based [60], Difference-based [26], Interlaced-based [46], and Combine-based methods [18, 47]. As mentioned above, many open-source methods have a narrow focus on their specific models, making it challenging to extend to multiple frameworks. However, OpenStereo provides a solution to this problem by supporting all of the aforementioned frameworks. With OpenStereo, researchers and practitioners can easily compare and evaluate different stereo matching models under a standardized evaluation protocol.
Support for various evaluation datasets. OpenStereo is a comprehensive tool that not only supports synthetic stereo datasets such as SceneFlow [35], but also five real-world datasets: KITTI2012 [14], KITTI2015 [36], ETH3D [45], Middlebury [43], and DrivingStereo [66]. We introduce a suite of bespoke functions, meticulously crafted for each dataset, encompassing everything from initial data preprocessing to the final stages of evaluation. The evaluation module includes the submission of the results to KITTI2012 [14] and KITTI2015 [36] leadboards.
@@
3.2 Main Modules
Technically, we follow the design of most PyTorch deep learning projects and divide OpenStereo into three modules, data, modeling, and evaluation.
Data module includes three key components: data loader, which is responsible for loading and preprocessing the data; data sampler, which is used to select a subset of data for each iteration; and data transform, which applies various transformations to the data to increase the diversity and complexity of the training set.
Modeling module is built upon the BaseModel base class, which predefines many behaviors of the stereo matching during both training and testing phases. This module consists of four essential components, namely the Feature Extraction, Cost Calculation, Cost Aggregation, and Disparity Processor, which are critical to current stereo matching algorithms.
Evaluation module is utilized to assess the performance of the obtained model. Recognizing that different datasets require distinct evaluation protocols, we have incorporated these varied protocols directly into OpenStereo, significantly simplifying the evaluation process.
@@ 4 Revisit Deep Stereo Matching
4.1 Recheck on Previous Methods
Figure 3: Quantitative evaluation on the SceneFlow [35] test and KITTI2015 [36] leadboard. For each model, the specific category on the SceneFlow used is consistent with the original paper. Underline refers to evaluation in the non-occluded regions only STTR [31].
For benchmarking, it is critical to ensure that the results are reliable and trustworthy. To achieve this, we conducted our experiments on SceneFlow [35] and KITTI2015 [36] datasets. As shown in Figure 3, the reproduced performances of OpenStereo are better than the results reported by the original papers. Regarding the KITTI2015 dataset, submission constraints led us to limit our leaderboard contributions to reproductions of the widely recognized PSMNet [6] and the latest state-of-the-art IGEV [60]. OpenStereo is designed to offer the research community in stereo matching a standardized, comprehensive platform for method assessment. This facility enables meaningful and comparative analyses across various models, thereby serving as an valuable tool for the innovation and scrutiny of emerging algorithms.
4.2 Denoising Stereo Matching
Figure 4: The visualization of stereo images and disparity with different data augmentation. The blue box represents the area where random cropping occurs during training. Notably, when the views are horizontally flipped, the disparity are multiplied by 11-1- 1, making the disparity map appear pure white.
Table 1: The model undergoes training and evaluation on the finalpass of the SceneFlow [35] training and test sets. KITTI2015 [36] training set, consisting of 200 images, is only employed to evaluate the generalizability of models. RC stands for RandomCrop. HFlip denotes both images of a stereo and disparity are horizontally flipped. HSFlip horizontally flips both images in the stereo pair and then swaps them. VFlip involves vertically flipping both images in the stereo pair along with the disparity, inverting their top-bottom orientation. CES represents ColorAug, Erase, and Scale. Settings used in our final model are underlined.
@@ 4.2 Denoising Stereo Matching
Table 2: Ablation study of different backbones on the finalpass of SceneFlow test datasets [35]. Flops and Params represent the computational complexity and parameters within the whole model, respectively. FLOPs are calculated at a resolution of $544\times 960544 × 960$.

### v7

@@ Abstract
Stereo matching aims to estimate the disparity between matching pixels in a stereo image pair, which is important to robotics, autonomous driving, and other computer vision tasks. Despite the development of numerous impressive methods in recent years, determining the most suitable architecture for practical application remains challenging. To address this gap, our paper introduces a comprehensive benchmark focusing on practical applicability rather than solely on individual models for optimized performance. Specifically, we develop a flexible and efficient stereo matching codebase, called OpenStereo. OpenStereo includes training and inference codes of more than 10 network models, making it, to our knowledge, the most complete stereo matching toolbox available. Based on OpenStereo, we conducted experiments and have achieved or surpassed the performance metrics reported in the original paper. Additionally, we conduct an exhaustive analysis and deconstruction of recent developments in stereo matching through comprehensive ablative experiments. These investigations inspired the creation of StereoBase, a strong baseline model. Our StereoBase ranks 1st on SceneFlow, KITTI 2015, 2012 (Reflective) among published methods and performs best across all metrics. In addition, StereoBase has strong cross-dataset generalization. Code is available at https://github.com/XiandaGuo/OpenStereo.
@@ 1 Introduction
Traditional stereo-matching algorithms typically match corresponding image regions between the left and right views based on their similarity measures. Several techniques have been proposed in the literature for stereo matching, including methods based on gray-level information [3, 26, 62], region-based approaches [67, 34], and energy optimization methods [38, 18]. With the support of large synthetic datasets [30, 39, 40, 12, 31, 61], CNN-based stereo matching methods [21, 5, 15, 55] has achieved impressive results. As shown in Figure 1, based on the network pipeline of stereo matching, CNN-based stereo matching methods can be roughly grouped into two categories [48], including the encoder-decoder network with 2D convolution (ED-Conv2D) [30, 35, 60, 36, 48, 56, 45, 28, 29, 25, 51, 27] and the cost volume matching with 3D convolution (CVM-Conv3D) [21, 5, 64, 15, 65, 10, 68, 13, 1, 8, 2, 42, 54, 55, 7, 57].
However, we find that different studies often employ various data augmentation strategies, learning rates, learning rate optimization methods, and backbone architectures. This inconsistency makes it difficult to evaluate and compare various methods’ performance accurately. This inconsistency in experimental setups and methodologies makes it difficult to derive conclusive insights and hampers the objective assessment of advancements in stereo matching. Without a standardized benchmark, researchers struggle to identify the true impact of individual components and innovations. There is a lack of clear conclusions and exploration regarding data augmentation strategies, backbone selection, and cost construction methods in stereo matching.
Moreover, not all datasets are accompanied by official evaluation tools. The SceneFlow [30] dataset, with its finalpass and cleanpass data varieties, complicates fair model comparisons. Generalization experiments for stereo matching algorithms typically train on the SceneFlow dataset and evaluate on KITTI2012 [12], KITTI2015 [31], ETH3D [40], and Middlebury [39]. However, due to the absence of a standard protocol for generalization experiments, different papers may yield inconsistent results for the same method. For instance, discussions on the generalization performance of IGEV [55] across works [55, 49, 14] exemplify this issue.
@@
Hence, there’s a pressing need for a comprehensive benchmark study within the stereo-matching community to enhance practicality and ensure consistent comparisons. To achieve this objective, we introduce a versatile stereo-matching codebase: OpenStereo. To promote scalability and adaptability, OpenStereo offers the following features: (1) Modulal design, researchers can define a new model without the need to alter the model code itself by simply modifying a YAML configuration file. (2)Various frameworks, including Concatenation-based [5, 64], Correlation-based [56, 47, 48, 2], Interlaced-based [41], Group-wise-correlation-based [55], Combine-based methods [15], and Difference-based [22]. (3)Various datasets, including SceneFlow [30], KITTI2012 [12], KITTI2015 [31], Middlebury [39], and ETH3D [40]. (4) State-of-the-art methods, including PSMNet [5], GwcNet [15], AANet [56], FADNet++ [48], CFNet [42], STTR [28], CoEx [2], CascadeStereo [13], MobileStereoNet [41] and IGEV [55].
Leveraging OpenStereo, we rigorously reassess various officially stated conclusions by re-implementing the ablation studies, including data augmentation, backbone architectures, cost construction, disparity regression, and refinement processes. Based on the insights gleaned from these ablation experiments, we introduce StereoBase, a model that sets a new benchmark, surpassing recently proposed methods in terms of performance. StereoBase is powerful and serves as an empirically state-of-the-art (SOTA) baseline model for stereo matching, demonstrating exceptional efficacy and resilience across diverse testing scenarios. Our contributions are summarized as follows:
@@
  • $\bullet$
    We evaluate the impact of CNN, Transformer, and hybrid CNN-Transformer backbones on stereo matching. To our knowledge, this is the first work to explore these architectures.
@@ 2.1 Stereo Matching
Figure 2: The design principles of proposed codebase OpenStereo.
Stereo Matching with CVM-Conv3D The CVM-Conv3D methods are proposed to improve the performance of depth estimation [21, 5, 59, 50, 64, 15, 65, 10, 68, 13, 1, 8, 2, 42, 54, 55, 57]. These methods learn disparities from a 4D cost volume, mainly constructed by concatenating left feature maps with their corresponding right counterparts across each disparity level [5]. GCNet [21] firstly introduced a novel approach that combines 3D encoder-decoder architecture with a 2D convolutional network to obtain a dense feature representation, which is used to regularize a 4D concatenation volume. Following GCNet [21], PSMNet [5] proposes an approach for regularizing the concatenation volume by leveraging a stacked hourglass 3D convolutional neural network in tandem with intermediate supervision. To enhance the expressiveness of the cost volume and ultimately improve performance in ambiguous regions, GwcNet [64] proposes the group-wise correlation volume and ACVNet [54] proposes the attention concatenation volume. CoEx [2] proposes a novel approach called Guided Cost volume Excitation (GCE), which leverages image guidance to construct a simple channel excitation of the cost volume. IGEV-Stereo [55] leverages an iterative geometry encoding volume to capture local and non-local geometry information, outperforming existing methods on KITTI benchmarks and achieving cross-dataset generalization and high inference efficiency.
@@
Stereo Matching with ED-Conv2D The ED-Conv2D methods [30, 35, 60, 36, 47, 48, 56, 45, 28, 29, 41, 25, 51, 27, 16], which adopt networks with 2D convolutions to predict disparity, has been driven by the need for improved accuracy, computational efficiency, and real-time performance. One of the early deep learning-based stereo matching methods, MC-CNN (Matching Cost CNN) [63], was proposed to learn a matching cost function for improving performance in the cost aggregation and optimization stages. Subsequently, Mayer et al [30] present end-to-end networks for the estimation of disparity, called DispNet, which is pure 2D CNN architectures. However, the model still faces challenges in capturing the matching features, resulting in poor estimation results. To overcome this challenge, the correlation layer is introduced in the end-to-end architecture [30, 9, 19, 20] to better capture the relationship between the left and right images. By incorporating this layer, the accuracy of the model is significantly improved. Furthermore, FADNet++ [48] proposes an innovative approach to efficient disparity refinement using residual learning [17] in a coarse-to-fine manner. AutoDispNet [36] applied neural architecture search to automatically design stereo-matching network structures. More recently, Croco-Stereo [51] shows that large-scale pre-training can be successful for stereo matching through well-adapted pretext tasks. This method can achieve state-of-the-art performance without using task-specific designs, like correlation volume or iterative estimation.
@@ 3 OpenStereo
In recent years, there has been a proliferation of new frameworks and evaluation datasets for stereo-matching. However, the lack of a unified and fair evaluation platform in this field is a significant issue that cannot be ignored. To address this challenge and promote academic research and practical application we have developed OpenStereo, a pyTorch-based [33] toolbox that provides a reliable and standardized evaluation framework for stereo matching.
@@ 3.1 Design Principles of OpenStereo
Compatibility with various frameworks. Currently, more and more stereo matching methods have emerged, such as Concatenation-based [21, 5, 64, 8], Correlation-based [56, 47, 48, 2], Group-wise-correlation-based [55], Difference-based [22], Interlaced-based [41], and Combine-based methods [15, 42]. As mentioned above, many open-source methods have a narrow focus on their specific models, making it challenging to extend to multiple frameworks. However, OpenStereo provides a solution to this problem by supporting all of the aforementioned frameworks. With OpenStereo, researchers and practitioners can easily compare and evaluate different stereo-matching models under a standardized evaluation protocol.
Support for various evaluation datasets. OpenStereo is a comprehensive tool that not only supports synthetic stereo datasets such as SceneFlow [30], but also four real-world datasets: KITTI2012 [12], KITTI2015 [31], ETH3D [40], and Middlebury [39]. We introduce a suite of bespoke functions, meticulously crafted for each dataset, encompassing everything from initial data preprocessing to the final stages of evaluation. The evaluation module includes the submission of the results to KITTI2012 [12] and KITTI2015 [31] leadboards.
@@ 4 Revisit Deep Stereo Matching
Figure 3: Quantitative evaluation on the SceneFlow [30] and KITTI2015 [31] leadboard. For each model, the specific category on the SceneFlow used is consistent with the original paper. Underline refers to evaluation in the non-occluded regions only STTR [28].
4.1 Datasets and Evaluation Metrics
SceneFlow [30] is a synthetic stereo collection, counting 35,454/4,370 image pairs for training/testing, respectively. The data is split into two categories: cleanpass and finalpass. Cleanpass refers to the synthetic images that are generated by clean renderings without post-processing, while finalpass images are produced with photorealistic settings such as motion blur, defocus blur, and noise. The end point error (EPE) is utilized as the evaluation metric.
KITTI. KITTI 2012 [12] and KITTI 2015 [31] are datasets that captured from real-world scenes. KITTI 2012 contains 194/195 image pairs for training/testing, while KITTI 2015 provides 200/200 image pairs for training and testing. All KITTI datasets provide sparse ground-truth disparities guided by the LiDAR system. For evaluations, we calculate EPE and the percentage of pixels with EPE larger than 3 pixels in all (D1-all) regions. All two KITTI datasets are also used for cross-domain generalization performance evaluation.
Middlebury 2014 [39] has two sets of 15 image pairs for both training and testing, respectively. The stereo images are captured from indoor scenes, with 3 resolution options, which are full, half, and quarter. In evaluations, only the training image pairs with half resolution will be used to evaluate the cross-domain generalization performance, and EPE and 2px Error metric are reported.
ETH3D [40] has stereo images collected from both indoor and outdoor environments with grayscale formate. It comes with 27 and 20 image pairs for training and testing, respectively, and a sparsely labeled ground truth for reference. In evaluation, only the training set is used to report the cross-domain generalization performance. And we adopt the EPE and 1px Error as the metric.
4.2 Evaluation of Prior Work
For benchmarking, it is critical to ensure that the results are reliable and trustworthy. To achieve this, we conducted our experiments on SceneFlow [30] and KITTI2015 [31] datasets. Regarding the KITTI2015 dataset, submission constraints led us to limit our leaderboard contributions to reproductions of the widely recognized PSMNet [5] and the latest SOTA IGEV [55]. (More details in the Supplementary Material). As shown in Figure 3, the reproduced performances of OpenStereo are better than the results reported by the original papers. This is primarily due to the normalization applied to the input stereo images by the OpenStereo framework. (All other settings remain consistent with the original paper.)
4.3 Necessity of Comprehensive Ablation Study
In the evolving landscape of stereo-matching, comprehensive ablation studies play a pivotal role in deciphering the effectiveness of different components and strategies. A thorough ablation study goes beyond mere performance metrics; it uncovers the underlying mechanics of different algorithms, revealing their strengths and weaknesses in various scenarios. For instance, different data augmentation techniques may yield contrasting effects on the model’s ability to match stereo images accurately. Similarly, the impact of various backbones, cost volume configurations, and disparity regression methods on the overall performance can be profound. Understanding the specific contributions of each component is crucial for building more efficient and effective stereo-matching systems.
@@ 4.3 Necessity of Comprehensive Ablation Study
Table 1: Ablation study on Data Augmentation and LR_scheduler Selection. KITTI2015 [31] training set, consisting of 200 images, is only employed to evaluate the generalizability of models. RC stands for RandomCrop [23]. HFlip [23] denotes both images of a stereo and disparity are horizontally flipped. HSFlip [23] horizontally flips both images in the stereo pair and then swaps them. VFlip [23] involves vertically flipping both images in the stereo pair along with the disparity, inverting their top-bottom orientation. CES represents ColorAug [23], Erase [69], and Scale [43]. Settings used in our final model are underlined.
4.4 Denoising Stereo Matching
ImpletionDetails For the experiments of this subsection, we adopt the COEX [2] as the baseline model due to its speed. All experiments in this subsection are conducted on SceneFlow [30] datasets. FLOPs are calculated at a resolution of $544\times 960544 × 960$. The total epochs is set to 50.
Figure 4: The visualization of stereo images and disparity with different data augmentation. The blue box represents the area where random cropping occurs during training. Notably, when the views are horizontally flipped, the disparity is multiplied by 11-1- 1, making the disparity map appear pure white.
4.4.1 LR_scheduler and Data Augmentation
As shown in Table 1, MultiStepLR yields an EPE of 0.6839, while OneCycleLR achieves a lower EPE of 0.6155. This substantial difference underscores the crucial role of selecting an appropriate learning rate scheduler for stereo matching. The superior performance of OneCycleLR indicates its potential to improve model accuracy and robustness, making it a preferable choice over MultiStepLR for training stereo-matching models. Although five data augmentation techniques—random crop, color augmentation, eraser transform, flip, and spatial transform—are commonly used in stereo matching [29, 55], their empirical efficacy specifically for stereo matching has not been thoroughly explored. This study investigated these data augmentation strategies to address this gap. Visualizations of these augmentations are provided in Figure 4 for further insight. Most data augmentation strategies, except for the combined use of HSFlip and random crop, lead to a decline in the model’s EPE metric of SceneFlow. This is because stereo matching involves pixel-level matching, and these data augmentations (color augmentation and spatial transform) can affect the alignment of pixels. The combination of ColorAug, Erase, and Scale (CES) shows the best generalization performance on KITTI2015, with the lowest EPE of 1.56 and D1_all of 7.64, although it increases the EPE on SceneFlow to 0.7240. These findings underscore the importance of selecting appropriate data augmentation methods to enhance model accuracy and robustness.